FEL Upgrade Project Weekly Brief
September 10-14, 2001

Special Message:
The members of the FEL team offer our condolences to our friends and colleagues who lost loved ones in this week’s tragic
events in NYC, the Pentagon and the US airlines.  We realize that this week’s events will change how all of us in the civilized
world go about our daily lives.  We also realize that for all of our professional lives, as we worked in the science and
technology community, our work, our basic and applied science and technology development has been supported directly by
the American people through our federal, state and local governments.   We are extremely grateful for this support.  We stand
ready to continue to use our brains and hands for our national needs where our experience can best be used.

It has been a difficult and emotional week for all of us.  Our work continues despite the obvious interruptions this week for
condolences and concerns about security.  Our report continues.

Despite having our RF engineers stranded across the country by the shutdown of the airlines this week, CPI in Palo Alto
proceeded with the scheduled test of 100 kW klystrons for the upgrade injector.  CPI is pleased to report to us that the tube
meets all specifications and has delivered 112 kW.

We held a preliminary design review on the optical cavity design for the FEL upgrade today,
(Sept. 14th)

We received funding ($2.84M) from the Air Force for the start up of the 1 kW UV FEL program this week in the remaining
weeks of FY01.  Fortunately, the account is valid until Sept. 30, 2002.

In addition, we have received information that additional funding for the project is included in
both the SASC and HASC FY02 authorization bills for the Dept. of Defense.

We are pleased to report that the Jefferson Lab DOE Site Office received on Sept. 11th the funding and authorization from
the Air Force Research Lab for JLab to proceed with the first phase of the UV FEL project in collaboration with Aerospace
Corp.  Jlab received authorization from DOE to proceed with the project on Sept.13th .  We thank John Eric at AFRL and
Wayne Skinner, our DOE contracting officer for their assistance with the final paper work.

Our financial data was received for the month of August so we proceeded with our analysis of the close-out of the Phase 1
contract by Sept. 30.

We held a safety inspection this week and had only minor issues to deal with.  They were being cleaned up on Friday.  One
previous safety issue regarding possible high N2 flow into the Optical Clean Room was corrected by the addition of a flow

Three proposals (Cryogenic Mirror Development, Injector Drive Laser Development, and Beam Position Monitor
Development) for the new JTO submission were written and circulated around JLab for approval.  Preparations continued for
the DOE/SURA review of JLab’s “Science and Technology” which is scheduled for 9/24-26.

WBS 3 (Beam Physics):
Vacuum systems were discussed. Most effort went toward refinement and documentation of
the CEBAF-ER design, a proposed high energy test of the energy recovery concept.

WBS 4 (Injector):
Work continued on the cathode shield assembly. The cathode stalk subassembly was brazed and is back for final machining.
The cathode ball was roughed out inside and is having the interior electro-polished prior to final machining. Machining on the
gun chamber main body was started. Assisted W&M implantation efforts by cleaning their modified chamber.

Completed the construction of the first production 7-cell cavity.

WBS 6 (RF):
As noted in the highlights, we received notification from CPI that the 100 kW klystron has met or exceeded all JLab
specifications during the scheduled acceptance tests this week.

WBS 8 (Instrumentation):
Prototype boards for a 4 channel combination amplifier/line driver have received and have successfully completed their initial
testing.  A design for the TTL version has been completed and prototype boards will be in soon.

The video system, channel 16, is available in various formats at http://laser.jlab.org/video.

Beam Position Monitors and Optical Beam Position Monitors underwent further testing using the FEL.  Results to follow.

Multiple cables are now available for all users in the user labs.  These include various lengths of CAT 5 cable and BNC

Drawings completed this week:  OBPM Molectron Quad Buffer Schematic, Fabrication, Assembly and Artwork; OBPM
Molectron Quad Daughter Board Schematic, Fabrication, Assembly and Artwork; Beamviewer Interlock and Lamp/Camera
Interface Box Fabrication and Assembly; OBPM Circuit Board Test Box Wiring Diagram and Fabrication Details; and the 32
Channel Sample & Hold Buffer Board Schematic, Assembly, Fabrication and Artwork.

WBS 9 (Transport):
Optical Chicane Dipoles (GW)
o  Clarifications were received on the GW bids and forwarded to G. Biallas for review.
    Expecting BAFO next week.
o  The copper conductor for the GW's was received and the steel is being shipped.
Injector Dipoles (GU/GV)
o  The final round of rework on the Small Magnet's (GV) Assembly Drawing is nearly complete.  They started the final
correction of the Large Magnet’s (GU) details.
Arc Dipoles (GY, GX, GQ)
o  AES started adjusting the DY drafting model (and hence, automatically, most of the details) to the dimensions of the
successful magnetic model.
3 inch quad (QX)
o  Work continues on cores and coils.  Request for bids on assembly went out on Sept. 7.  Bids due Sept 26.
Trim Quad (QT)
o  In procurement.
Sextupole (SF)
o  Draft drawings have been started.  Finally back to work on the 3D model.
Octupole (OT)
o  DULY Research adjusted the pole tip shape to counter the end effects from this "all ends" magnet and are well within the
required 1% quality.  They added a field clamp to reduce the stray field and the quality went out of specification.  The are
working on shaping the field clamp along with further adjustments to the pole tip to bring the quality within specification.
Beam Line and Vacuum
o  Work continued on the definition of the Optical Chicane chambers and continued to the mounting stand for the GW
o  The bid package drawing set for the girders in the return leg continues in detailing.

WBS 10 (Wiggler):
The wiggler measurement apparatus was worked on this week to improve its stability and fix some bugs in the apparatus.  We
expect to begin production scans next week.

One wiggler chamber was welded this week and sent out for final machining.  The other is being cleaned for welding.  The first
chamber is expected to be complete next week.  The final chamber will be done the following week.

WBS 11 (Optics):
Drawings of some custom components for the optical cavity were signed off and procurement is in process.

The design of the cooling loop and electrical connections is in the conceptual stage, detailing continues in other areas.

We prepared for, and will hold, an internal review of the optical cavity design today (9/14).

We have set up the experimental hardware to do the next set of material exposures funded by the JTO contract.

Preparations continued for the JTO damage studies to be resumed next week.  NRL folks are here setting up their
equipment.  A brief test run with the FEL was made for the nanotube collaboration so that they could test their new exposure
chamber which has cross ports for diagnostic access.